OpenStudy (anonymous):

simplify [(-1)^5/(-2)^-3]^2

OpenStudy (anonymous):

(-1)^5 = -1 (-2)^-3 = -1/8 so your problem can be simplified to: [ -1/(-1/8) ]^2 = 8

OpenStudy (anonymous):

that's wrong... it should be 64